OpenStudy (anonymous):

Find the equation of the line tangent to y^2 = -16x and parallel to x+y=1

OpenStudy (anonymous):

@satellite73

OpenStudy (tkhunny):

Please find the slope of the line that sets the standard of parallel.

OpenStudy (anonymous):

y=-x+1 so the slope is -1

OpenStudy (tkhunny):

Good. Can you find a general expression for dy/dx from the parabola?

OpenStudy (anonymous):

wait, we are not supposed to use calculus. only analytic geometry expressions are allowed. :)

OpenStudy (tkhunny):

I wish I knew how to tell what section posts were in. You haven't a plan for that? Did you first get a picture?

OpenStudy (anonymous):

i have seen these before, the non calculus version you have get a quadratic equation and assert that there is only one solution, set the discriminant equal to zero and solve it is a pain lets see if i can remember it

OpenStudy (anonymous):

nope, but i have the answer, here it is: x + y=4

OpenStudy (anonymous):

@satellite73 yeah, it is the right procedure

OpenStudy (anonymous):

@tkhunny have you seen this? i have seen it here a couple times, always forget, but maybe we can recreate it

OpenStudy (anonymous):

oh yeah it is not that bad i think

OpenStudy (tkhunny):

Workign on it. It's only quadratic, so it shouldn't be too bad. They were doing this kind of thing long before the calculus was invented.

OpenStudy (anonymous):

\[y^2=-16x\\ y=-x+b\] set them equal \[(-x+b)^2=-16x\] and you want only one solution to this, solve for \(b\)

OpenStudy (anonymous):

i better do it with pencil an paper first but the idea is that \[b^2-2bx+16x+b^2=0\] should have only one solution \[x^2+(16-2b)x+b^2=0\] set the discriminant equal to zero and solve for \(b\)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

i already solve b=4

OpenStudy (anonymous):

ok then you beat me to it

OpenStudy (anonymous):

so y=-x + b y=-x + 4 x+y=4 alright. thanks to you @satellite73 and @tkhunny for helping! :)

OpenStudy (anonymous):

wait. another question for the two of you. find the equations of the lines tangent to y^2= 4x containing (-2,1)

OpenStudy (anonymous):

@satellite73

OpenStudy (tkhunny):

No calculus rule, again?

OpenStudy (anonymous):

yup

OpenStudy (tkhunny):

We should be able to get this far without much trouble. Pick an arbitrary point on the parabola, (a,b), and write the equation of the line associated with it. \(y-1 = \dfrac{b-1}{a+2}(x+2)\) Enforcing the relationship between a and b, we get. \(y-1 = \dfrac{b-1}{\dfrac{b^{2}}{4}+2}(x+2)\) or \(y-1 = 4\cdot\dfrac{b-1}{b^{2}+8}(x+2)\) Did you get that far?

OpenStudy (anonymous):

yeah. what should i do next?

OpenStudy (anonymous):

equate it to y^2=4x

OpenStudy (anonymous):

equate it to y^2=4x?

OpenStudy (tkhunny):

Then you apply the old geometer's trick. If the point on the top is \((a_{1},b_{1})\), the equation of that tangent is \(b_{1}y = 2(a_{1} + x)\) If the point on the bottom is \((a_{2},b_{2})\), the equation of that tangent is \(b_{2}y = 2(a_{2} + x)\) Since both have to pass through (-2,1), this gives the relationship between the two a and b.

OpenStudy (tkhunny):

In other words, for the top one, the slope is \(\dfrac{2}{b_{1}}\) and for the bottom one, the the slope is \(\dfrac{2}{b_{2}}\) and this completes the exercise.